# Scale-Invariant Gradient Aggregation for Constrained Multi-Objective Reinforcement Learning

### v1

@@ Abstract
Multi-objective reinforcement learning (MORL) aims to find a set of Pareto optimal policies to cover various preferences. However, to apply MORL in real-world applications, it is important to find policies that are not only Pareto optimal but also satisfy pre-defined constraints for safety. To this end, we propose a constrained MORL (CMORL) algorithm called Constrained Multi-Objective Gradient Aggregator (CoMOGA). Recognizing the difficulty of handling multiple objectives and constraints concurrently, CoMOGA relaxes the original CMORL problem into a constrained optimization problem by transforming the objectives into additional constraints. This novel transformation process ensures that the converted constraints are invariant to the objective scales while having the same effect as the original objectives. We show that the proposed method converges to a local Pareto optimal policy while satisfying the predefined constraints. Empirical evaluations across various tasks show that the proposed method outperforms other baselines by consistently meeting constraints and demonstrating invariance to the objective scales.
Reinforcement Learning, Multi-Objective, Deep RL
@@ 1 Introduction
Many real-world reinforcement learning (RL) applications involve multiple objectives. For example, in autonomous driving, vehicles must maintain fuel, avoid collisions, and reach their destination quickly. The presence of such multiple objectives has led to extensive studies of multi-objective RL (MORL) (Roijers et al., 2013), which aim to find a set of Pareto-optimal policies (Hayes et al., 2022) that are not dominated by any other policies, instead of finding a single optimal policy. This allows users to choose and utilize their preferred policy from the set without additional training. In order to represent the set of Pareto optimal policies, it is common to use a concept called preference (Yang et al., 2019; Kyriakis & Deshmukh, 2022), which encodes the relative importance of the objectives. A preference-conditioned policy can then represent the set of policies.
A number of MORL algorithms (Désidéri, 2012; Yang et al., 2019; Basaklar et al., 2023) update the preference-conditioned policy by maximizing a scalarized reward, computed as the dot product of a given preference and rewards from the multiple objectives. However, such scalarization is sensitive to the scale of objectives. For example, if a particular reward among multiple rewards has a significant scale, the majority of the trained policy set will converge towards focusing on that reward, as observed in Section 6.5. Hence, despite training a set of policies, there may be no user-preferred policy in the set, as they tend to be clustered around high-scale rewards. For this reason, it is required to develop a scale-invariant MORL algorithm.
Figure 1: An example of constrained MORL (CMORL). The robot aims to maximize energy efficiency and velocity while maintaining its balance to avoid falling. To this end, CMORL finds a set of feasible policies that are not dominated by other policies and satisfy constraints, which are indicated as the dashed line.
@@
Additionally, in domains where safety is paramount, such as robotic control and autonomous driving, it is crucial to incorporate constraints into MORL to enhance safety assurance, an example of which is shown in Figure 1. A straightforward approach to reflect constraints in MORL is treating the constraints as other objectives and concurrently updating the preferences corresponding to those objectives to maintain them below thresholds, as done by (Huang et al., 2022). This approach exactly aligns with the Lagrangian method, which solves the Lagrange dual problem by updating policy and multipliers alternatively. However, the Lagrangian method can make training unstable due to the concurrent update of policy and multipliers (Stooke et al., 2020; Kim & Oh, 2022). As the updates of the policy and multipliers influence each other, any misstep in either can make the training process diverge. Therefore, it is essential to incorporate constraints into MORL without additional optimization variables.
In this paper, we propose a constrained MORL (CMORL) algorithm that is scale-invariant and does not require any additional optimization variables. The primary component of our algorithm is Constrained Multi-Objective Gradient Aggregator (CoMOGA), which calculates a policy gradient by aggregating the gradients of the objective and constraint functions. Instead of treating the constraints as new objectives like an existing CMORL method (Huang et al., 2022), CoMOGA transforms the objectives into extra constraints. This novel transformation process makes solving the CMORL problem simple and intuitive by removing the multi-objective components. The transformed constraints are designed to increase the objectives using the gradients of the objective functions. Since only the directional component of the gradient is used, they are invariant to the objective scales. Furthermore, to reflect a given preference, we set the threshold of the transformed constraints to be proportional to the preference value. CoMOGA finally calculates the policy gradient by solving the transformed problem using linear approximation within a local region of the policy parameter space. Under mild assumptions, we have shown that updating a policy using CoMOGA monotonically increases objectives corresponding to specific preferences while consistently satisfying the constraints. Thus, constraints can be handled using CoMOGA alone, eliminating the need for additional variables. Furthermore, we analyze the convergence of the proposed method from a Pareto-optimal perspective and show that the proposed method converges to a local Pareto-optimal policy.
In order to evaluate the proposed method, experiments are conducted across diverse environments, including multi-objective tasks with and without constraints. We use well-known MORL metrics, namely hypervolume and sparsity (Xu et al., 2020), to gauge performance. The experimental results demonstrate that the proposed method achieves the best hypervolume and sparsity in the majority of tasks while successfully adhering to the constraints in all tasks. Finally, our contributions are summarized as follows:
  • We propose a scale-invariant CMORL algorithm that does not require additional optimization variables.
  • We theoretically show that the proposed method keeps satisfying constraints and converges to a local Pareto optimum under mild assumptions.
  • The proposed method outperforms baseline methods in terms of MORL metrics and constraint satisfaction in various experiments.
@@ 2 Related Work
Multi-Objective RL. MORL aims to find a set of Pareto-optimal policies that are not dominated by each other. Given that Pareto-optimal policies can be obtained with appropriate reward scalarization (Hayes et al., 2022), many MORL algorithms utilize scalarization techniques. Based on linear scalarization, Yang et al. (2019) introduced a new Bellman operator for MORL algorithms, which estimates a target Q value for a given preference by combining Q values from other preferences. Basaklar et al. (2023) presented a similar Bellman operator and extended it to an actor-critic framework to enable its application in continuous action spaces. For the continuous action spaces, Lu et al. (2023) proposed a SAC-based MORL algorithm, theoretically demonstrating that elements of the Pareto-optimal policy set can be achieved through linear scalarization. Not only the linear scalarization, Van Moffaert et al. (2013) proposed a nonlinear scalarization method, called Chebyshev scalarization, which can cover a non-convex policy set. To further improve performance, the evolution strategy (ES) can be applied as done in multi-objective optimization criteria (Deb et al., 2000). Chen et al. (2020) and Xu et al. (2020) proposed methods that repetitively update a policy for each population using linear scalarization and produce a new generation of policy parameters. However, such scalarization exhibits sensitivity to the scale of objectives. To tackle this issue, Abdolmaleki et al. (2020) introduce a scale-invariant algorithm termed MOMPO, which does not use scalarization like our proposed approach. This is further extended to a CMORL algorithm (Huang et al., 2022). MOMPO finds a set of policies in the distribution space instead of the reward space by heuristically calculating non-parametric policies. Due to the heuristics, MOMPO does not ensure convergence to a local Pareto-optimal policy, but our method is not only scale-invariant but also guarantees convergence.
@@
Constrained RL. Constrained RL is developed to explicitly consider safety or constraints in RL. Based on the approach to handling constraints, constrained RL algorithms can be categorized into three types: primal-dual methods, primal methods, and dual methods. The primal-dual methods, also called Lagrangian methods, are designed to address Lagrange dual problems. Ding et al. (2022) introduced a natural policy gradient-based primal-dual method and demonstrated its convergence to an optimal policy at a specified convergence rate. Another primal-dual method, proposed by Bai et al. (2022), ensures that a trained policy results in zero constraint violations during evaluation. While these methods are straightforward to implement and can be integrated with standard RL algorithms, the training process can be unstable due to additional optimization variables (Lagrange multipliers) (Stooke et al., 2020). In contrast to the primal-dual method, the primal method addresses the constrained RL problem directly, so no additional optimization variables are required. Achiam et al. (2017) introduced a trust region-based primal method. This approach linearly approximates a constraint within a trust region and updates the policy through a line search. Xu et al. (2021) presented a natural policy gradient-based algorithm and demonstrated its convergence rate towards an optimal policy. In multi-constrained settings, Kim et al. (2023) proposed a primal method named SDAC to handle infeasible starting cases. Lastly, dual methods aim to resolve the constrained RL problem within the state distribution space rather than the policy space, and Lee et al. (2022) use this approach. However, because the dual methods still need to use Lagrange multipliers to handle constraints, we have decided to develop a CMORL algorithm based on the primal methods.
@@ 3 Background
Constrained Multi-Objective MDP. We define a constrained multi-objective Markov decision process (CMOMDP) as a tuple represented by $\langle S$, $A$, $\mathcal{P}$, $R$, $C$, $\rho$, $\gamma\rangle$, where $S$ is a state space, $A$ is an action space, $\mathcal{P}:S\times A\times S\mapsto[0,1]$ is a transition model, $R:S\times A\times S\mapsto\mathbb{R}^{N}$ is a vectorized reward function, $C:S\times A\times S\mapsto\mathbb{R}^{M}$ is a vectorized cost function, $\rho$ is an initial state distribution, and $\gamma$ is a discount factor. The reward function $R$ has an upper bound $R_{\mathrm{max}}$, satisfying $R_{\mathrm{max}}\geq|R_{i}(\cdot,\cdot,\cdot)|$ $\forall i$. In the CMOMDP setting, a policy $\pi(\cdot|s)\in\Pi$ can be defined as a state-conditional action distribution. The objective function of the $i$th reward is defined as:
$\small J_{R_{i}}(\pi):=\mathbb{E}\left[\sum\nolimits_{t=0}^{\infty}\gamma^{t}R_{i}(s_{t},a_{t},s_{t+1})\right],$ (1)
@@
$\mathrm{maximize}_{\pi}\;J_{R_{i}}(\pi)\;\forall i\in\{1,...,N\}$ (2)
$\;\mathrm{s.t.}\;J_{C_{k}}(\pi)\leq d_{k}\;\forall k\in\{1,...,M\},$
where $d_{k}$ is the threshold of the $k$th constraint. Here, it is difficult to determine which policy is optimal when there are multiple objectives. To address this, we instead define a set of optimal policies using the following notion, called constrained dominance (Miettinen, 1999):
@@
A policy $\pi$ which is not dominated by any policy is called a constrained-Pareto (CP) optimal policy, and the set of all CP optimal policies is called constrained-Pareto (CP) front (Miettinen, 1999). In conclusion, we can say that CMORL aims to find a subset of the CP front.
Preference. In order to express a subset of the CP front, many existing MORL works (Xu et al., 2020; Alegre et al., 2023) introduce a preference space, defined as: $\Omega:=\{\omega\in\mathbb{R}^{N}|\;\omega_{i}\geq 0,\;||\omega||_{\infty}=1\}$, [^1] It is common to define the preference space using the norm $||\omega||_{1}$, but for convenience, we employ $||\omega||_{\infty}$. Additionally, an one-to-one conversion exists between these definitions. where a preference $\omega\in\Omega$ serves to specify the relative significance of each objective. We denote the set of indices representing the maximum preference values as $I_{\omega}:=\{i|\omega_{i}=1\}$. Subsequently, a specific subset of the Pareto front can be characterized by a function mapping from the preference space to the policy space. We denote this function as a universal policy $\pi(\cdot|s,\omega)$, which is characterized as an action distribution contingent upon both the state and preference. As a result, our goal is to train the universal policy to cover the CP front as extensively as possible.
Linear Scalarization. Linear scalarization (LS) is a general approach to deal with the multi-objective optimization problem (2), and it is achieved by solving the following problem given a preference $\omega$:
[…]
The above can be addressed by converting it into a Lagrange dual problem and then concurrently updating the policy and Lagrange multipliers. An actor-critic-based MORL algorithm such as CAPQL (Lu et al., 2023) and PD-MORL (Basaklar et al., 2023) can be employed for the policy update. For more details, please refer to Appendix B.3.1. However, LS is sensitive to the scale of objectives. Intuitively, increasing the scale of the $i$th reward has the same effect as increasing $\omega_{i}$. Because of this, even if $\omega$ is the same, it may be mapped to a different policy depending on how the rewards are scaled. If training is conducted with an increased scale for the $i$th reward, the trained policy set may focus on maximizing the $i$th objective. As a result, LS is hard to guarantee that the policy set is evenly distributed in the reward space, which increases the likelihood that there is no user-preferred policy in the trained policy set. Therefore, we develop a scale-invariant CMORL algorithm without LS.
@@ 4 Proposed Method
In this section, our discussion will unfold in three parts: 1) introduce a method to calculate policy gradients for a single preference, 2) present a method for training a universal policy that encompasses a spectrum of preferences, and 3) conclude the section with an analysis of the convergence properties of the proposed method.
4.1 Constrained Multi-Objective Gradient Aggregator
We propose a gradient aggregation method called constrained multi-objective gradient aggregator (CoMOGA), which calculates a policy gradient to improve the objectives and comply with the constraints utilizing their respective gradients. The fundamental concept behind the proposed method is to transform the objectives into constraints to circumvent the difficulty in directly handling multiple objectives. For this transformation, we only use the directional component of the objective’s gradient, making the method invariant to the objective scales. We first introduce the transformation process in a simplified problem and extend it to the original CMORL problem. Then, the obtained constrained optimization problem is solved within a local region of the parameter space through linear approximation so as to avoid the need for additional optimization variables.
First, we parameterize a policy and define gradients of the objective and constraint functions, as well as a local region in the parameter space. By representing the policy as $\pi_{\theta}$ with a parameter $\theta\in\Theta$, the gradients of the objective and constraint functions in (2) are expressed as follows:
$\small g_{i}:=\nabla_{\theta}J_{R_{i}}(\pi_{\theta}),\;b_{k}:=\nabla_{\theta}J_{C_{k}}(\pi_{\theta}).$ (4)
We will denote the objectives as $J_{R_{i}}(\theta)$ and the constraints as $J_{C_{k}}(\theta)$ for brevity. The local region is then defined using a positive definite matrix $H$ for a general expression, as follows: $||\Delta\theta||_{H}\leq\epsilon$, where $||x||_{H}:=\sqrt{x^{T}Hx}$, the local region size $\epsilon$ is a hyperparameter, and $H$ can be the identity matrix or the Hessian of the KL divergence between the current and updated policies, as in TRPO (Schulman et al., 2015).
Figure 2: Process of CoMOGA. We visualize the process in the parameter space, and the gray areas represent constraints. (Linear approximation) CoMOGA linearly approximates the original CMORL problem in (2). The gradients of the objective and constraint functions are visualized as black and blue arrows, respectively. (Transformation) The objectives are converted to constraints as described in (8). The intersection of all constraints is shown as the red area. (Scaling) The solution of the transformed problem, $\bar{g}^{\mathrm{ag}}$, is scaled to ensure increasing the objectives and continuously satisfying the constraints.
Next, we introduce the process of converting objectives into constraints. We first present the transformation process in a simplified problem where only the $i$th objective exists and then extend it to the original problem. The problem of maximizing the $i$th objective within the local region is expressed as follows:
$\mathrm{max}_{g\in\Theta}\;J_{R_{i}}(\theta_{\mathrm{old}}+g)\quad\mathrm{s.t.}\;||g||_{H}\leq\epsilon.$ (5)

### v2

# Conflict-Averse Gradient Aggregation for Constrained Multi-Objective Reinforcement Learning

@@ Abstract
In many real-world applications, a reinforcement learning (RL) agent should consider multiple objectives and adhere to safety guidelines. To address these considerations, we propose a constrained multi-objective RL algorithm named Constrained Multi-Objective Gradient Aggregator (CoMOGA). In the field of multi-objective optimization, managing conflicts between the gradients of the multiple objectives is crucial to prevent policies from converging to local optima. It is also essential to efficiently handle safety constraints for stable training and constraint satisfaction. We address these challenges straightforwardly by treating the maximization of multiple objectives as a constrained optimization problem (COP), where the constraints are defined to improve the original objectives. Existing safety constraints are then integrated into the COP, and the policy is updated using a linear approximation, which ensures the avoidance of gradient conflicts. Despite its simplicity, CoMOGA guarantees optimal convergence in tabular settings. Through various experiments, we have confirmed that preventing gradient conflicts is critical, and the proposed method achieves constraint satisfaction across all tasks.
@@ 1 Introduction
Many real-world reinforcement learning (RL) applications involve multiple objectives and have to consider safety simultaneously. For instance, locomotion tasks for legged robots focus on maximizing goal-reaching success rate and energy efficiency while avoiding collisions with the environment, as illustrated in Figure 2. To effectively tackle these considerations, constrained multi-objective RL (CMORL) has been introduced [17]. CMORL aims to find a set of constrained-Pareto (CP) optimal policies [16], which are not dominated by any others while satisfying safety constraints, rather than finding a single optimal policy. This allows users to choose and utilize their preferred policy from the set without additional training. In order to express a set of CP optimal policies, it is common to use a concept called preference [34, 22], which encodes the relative importance of the objectives. Then, a set of policies can be represented by a preference-conditioned policy, a mapping function from a preference space to a policy space.
In most multi-objective RL (MORL) algorithms [12, 34, 7], the preference-conditioned policy is trained by maximizing a scalarized reward, computed as the dot product of a given preference and rewards of the multiple objectives. These approaches, called linear scalarization (LS), offer the benefits of straightforward implementation and ensure Pareto-optimal convergence in tabular settings [24]. However, when applied to deep RL with function approximators, LS tends to converge to local optima due to the nonlinearity of the objective functions, as shown in a toy example in Figure 2. In the field of multi-task learning (MTL) [35, 23, 27], this issue has been addressed by avoiding conflicts between the gradients of multiple objective functions. We hypothesize that similar improvements can be achieved in CMORL by avoiding gradient conflicts.
@@
It is also crucial to handle constraints as well as multiple objectives in CMORL. A straightforward approach is treating the constraints as other objectives and concurrently adjusting the preferences corresponding to those objectives to maintain them below thresholds, as done in [17]. This approach exactly aligns with the Lagrangian method, which solves the Lagrange dual problem by updating policy and multipliers alternatively. However, the Lagrangian method can make training unstable due to the concurrent update of policy and multipliers [30, 19]. As the updates of the policy and multipliers influence each other, any misstep in either can make the training process diverge. Thus, it is critical to manage constraints without introducing additional optimization variables.
In order to avoid gradient conflicts and efficiently handle safety constraints, we introduce a novel but straightforward algorithm named Constrained Multi-Objective Gradient Aggregator (CoMOGA). CoMOGA treats a multi-objective maximization problem as a constrained optimization problem (COP), where constraints are designed to enhance the original objective functions in proportion to their respective preference values. Existing safety constraints are then incorporated into the constraint set of the COP. By explicitly preventing the given objectives from decreasing, CoMOGA successfully avoids gradient conflicts. The policy gradient is then calculated by solving this transformed problem using linear approximation within a local region, thereby eliminating the need for extra variables to handle safety constraints. While the proposed method is intuitive and straightforward, we have also demonstrated that it converges to a CP optimal policy in a tabular setting.
The proposed method has been evaluated across diverse environments, including multi-objective tasks with and without constraints. The experimental results support our hypothesis that avoiding gradient conflicts is effective for preventing convergence to local optima. This is particularly essential in environments with unstable dynamics, such as bipedal and humanoid locomotion tasks, since they are prone to get stuck in suboptimal behaviors, such as frequent falling. Furthermore, the absence of additional variables enables stable training with constraint satisfaction across all tasks.
Figure 1: Example of CMORL. The robot aims to maximize energy efficiency and velocity while maintaining its balance to avoid falling. In order to consider such safety and multiple objectives, CMORL finds a set of feasible policies that are not dominated by other policies and satisfy constraints, which are indicated by the dashed line.
Figure 2: Comparison between LS and the proposed method. Optimization trajectories are indicated in red, with initial points marked by black circles. The contours illustrate the average values of the two objective functions, while the shaded area indicates regions where the constraints are violated. The grey line represents the CP optimal set. For some initial points, LS fails to reach the optimal set, whereas CoMOGA consistently finds it from any starting position. For more details, please see Appendix A.
@@ 2 Related Work
Multi-Objective RL. MORL aims to find a set of Pareto-optimal policies that are not dominated by any others. Given that Pareto-optimal policies can be obtained with appropriate reward scalarization [16], many MORL algorithms utilize scalarization techniques. Based on linear scalarization, Yang et al. [34] introduced a new Bellman operator for MORL algorithms, which estimates a target Q value for a given preference by combining Q values from other preferences. Basaklar et al. [7] presented a similar Bellman operator and extended it to an actor-critic framework to enable its application in continuous action spaces. For the continuous action spaces, Lu et al. [24] proposed a SAC-based MORL algorithm, theoretically demonstrating that elements of the Pareto-optimal policy set can be achieved through linear scalarization. Not only the linear scalarization, Van Moffaert et al. [31] proposed a nonlinear scalarization method, called Chebyshev scalarization, which can cover a non-convex policy set. To further improve performance, the evolution strategy (ES) can be applied as done in multi-objective optimization criteria [10]. Chen et al. [8] and Xu et al. [32] proposed methods that repetitively update a policy for each population using linear scalarization and produce a new generation of policy parameters. There is also a different approach from scalarization. Abdolmaleki et al. [1] have proposed a heuristic method of applying preference in the action distribution space instead of the reward space. It is further extended to a CMORL algorithm in [17].
@@
Constrained RL. Constrained RL is developed to explicitly consider safety or constraints in RL. Based on the approach to handling constraints, constrained RL algorithms can be categorized into two types: primal-dual methods and primal methods. The primal-dual methods, also called Lagrangian methods, are designed to address Lagrange dual problems. Ding et al. [11] introduced a natural policy gradient-based primal-dual method and demonstrated its convergence to an optimal policy at a specified convergence rate. Another primal-dual method, proposed by Bai et al. [6], ensures that a trained policy results in zero constraint violations during evaluation. While these methods are straightforward to implement and can be integrated with standard RL algorithms, the training process can be unstable due to additional optimization variables (Lagrange multipliers) [30]. In contrast to the primal-dual method, the primal method addresses the constrained RL problem directly, so no additional optimization variables are required. Achiam et al. [2] introduced a trust region-based primal method. This approach linearly approximates a constraint within a trust region and updates the policy through a line search. Xu et al. [33] presented a natural policy gradient-based algorithm and demonstrated its convergence rate towards an optimal policy. In multi-constrained settings, Kim et al. [20] proposed a primal method to handle infeasible starting cases.
@@ 3 Background
Constrained Multi-Objective MDP. We define a constrained multi-objective Markov decision process (CMOMDP) as a tuple represented by $\langle S$, $A$, $\mathcal{P}$, $R$, $C$, $\rho$, $\gamma\rangle$, where $S$ is a state space, $A$ is an action space, $\mathcal{P}:S\times A\times S\mapsto[0,1]$ is a transition model, $R:S\times A\times S\mapsto\mathbb{R}^{N}$ is a vectorized reward function, $C:S\times A\times S\mapsto\mathbb{R}^{M}$ is a vectorized cost function, $\rho$ is an initial state distribution, and $\gamma$ is a discount factor. The reward and cost function are bounded by $R_{\mathrm{max}}$. In the CMOMDP setting, a policy $\pi(\cdot|s)\in\Pi$ can be defined as a state-conditional action distribution. The objective function of the $i$th reward is defined as:
$J_{R_{i}}(\pi):=\mathbb{E}\left[\sum\nolimits_{t=0}^{\infty}\gamma^{t}R_{i}(s_{t},a_{t},s_{t+1})\right],$
@@
$\mathrm{maximize}_{\pi}\;J_{R_{i}}(\pi)\;\forall i\in\{1,...,N\}\quad\mathbf{s.t.}\;J_{C_{k}}(\pi)\leq d_{k}\;\forall k\in\{1,...,M\},$ (1)
where $d_{k}$ is the threshold of the $k$th constraint. Here, it is ambiguous to determine which policy is optimal when there are multiple objectives. To address this, we instead define a set of optimal policies using the following notion, called constrained dominance [25]:
@@
A policy $\pi$ which is not dominated by any policy is called a constrained-Pareto (CP) optimal policy, and the set of all CP optimal policies is called constrained-Pareto (CP) front [25]. Additionally, a condition for gradient conflicts is defined as follows:
Definition 3.2 (Gradient Conflict).
Let $g_{i}$ be the gradient of the $i$th objective function, and let $g$ be the gradient used for policy updates. We say there is a gradient conflict if $\exists i$ such that $g_{i}^{T}g<0$.
As studied in MTL [23, 35] and shown in Figure 2, gradient conflicts can cause convergence to local optima. Thus, we aim to find the CP front while updating policies without gradient conflicts.
Preference. In order to express a subset of the CP front, many existing MORL works [32, 5] introduce a preference space, defined as: $\Omega:=\{\omega\in\mathbb{R}^{N}|\;\omega_{i}\geq 0,\;||\omega||_{\infty}=1\}$, [^1] It is common to define the preference space using the norm $||\omega||_{1}$, but for convenience, we employ $||\omega||_{\infty}$. Additionally, an one-to-one conversion exists between these definitions. where a preference $\omega\in\Omega$ serves to specify the relative significance of each objective. Subsequently, a specific subset of the Pareto front can be characterized by a function mapping from the preference space to the policy space. We denote this function as a universal policy $\pi(\cdot|s,\omega)$, which is characterized as an action distribution contingent upon both the state and preference. Ultimately, our goal is to train the universal policy that covers the CP front as extensively as possible.
@@ 4 Proposed Method
In this section, our discussion is organized into four parts: 1) building a constrained optimization problem to handle the multiple objectives with safety constraints, 2) calculating a policy gradient by aggregating gradients of the objective and constraint functions for a single preference, 3) introducing a method for training a universal policy covering a spectrum of preferences, and 4) concluding with an analysis of the convergence properties of the proposed method. Before that, we first parameterize the policy and define gradients of the objective and constraint functions, as well as a local region in the parameter space. By representing the policy as $\pi_{\theta}$ with a parameter $\theta\in\Theta$, the gradients of the objective and constraint functions in (1) are expressed as:
$g_{i}:=\nabla_{\theta}J_{R_{i}}(\pi_{\theta}),\;b_{k}:=\nabla_{\theta}J_{C_{k}}(\pi_{\theta}).$
For brevity, we will denote the objectives as $J_{R_{i}}(\theta)$ and the constraints as $J_{C_{k}}(\theta)$. The local region is then defined using a positive definite matrix $H$ as: $||\Delta\theta||_{H}\leq\epsilon$, where $||x||_{H}:=\sqrt{x^{T}Hx}$, and $\epsilon$ is the local region size, a hyperparameter. $H$ can be the identity matrix or the Fisher information matrix, as in TRPO paper [29].
4.1 Transformation to Constrained Optimization
We introduce a process of treating the CMORL problem as a constrained optimization problem (COP). To provide insight, we first present the process in a simplified problem where only the $i$th objective exists and then extend it to the original problem. The problem of maximizing the $i$th objective within the local region is expressed as follows:
$\mathrm{max}_{\Delta\theta}\;J_{R_{i}}(\theta_{\mathrm{old}}+\Delta\theta)\quad\mathbf{s.t.}\;||\Delta\theta||_{H}\leq\epsilon.$ (2)